OpenStudy (i<3mybaby):

In the triangle below, calculate the sec θ. A. five-fourths B. five-thirds C. three-fourths D. four-fifths

OpenStudy (i<3mybaby):

@doulikepiecauseidont @iPwnBunnies

OpenStudy (arabpride):

hmm idk.... uhmmzz..... iz there a formula for sec??

OpenStudy (anonymous):

Yes, rememeber what we got for the formula of sec: 1/cos So if we find the cos of the angle we just have to put 1 over that

OpenStudy (i<3mybaby):

which is...

OpenStudy (anonymous):

Do you know the formulas for sin, cos and tan?

OpenStudy (anonymous):

\[\Large \sin =\frac{opp}{hyp}, \cos=\frac{adj}{hyp},\tan=\frac{opp}{adj}~or \frac{\sin}{\cos}\]

OpenStudy (i<3mybaby):

Not off the top of my head :(

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\Large \theta=Angle~measure \]

OpenStudy (anonymous):

So it's just a ratio (fraction) of the sides

OpenStudy (i<3mybaby):

WAIT WHAT IM CONFUSED

OpenStudy (anonymous):

Sorry I had to go, do you still need help with the question?

OpenStudy (i<3mybaby):

No I actually took and passed mymidterm today